# Opinion Clustering under the Friedkin-Johnsen Model: Agreement in Disagreement

### v2

@@ 5.1 Proof of Theorem-1
Proof 5.3.
By definition of $\alpha$, we get $\alpha^{c}\subset[n]$. From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$. If $\rho((I-\beta[\alpha^{c}])W[\alpha^{c}])<1$, then identity (5) follows by Neumann series. Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$. To prove that $\rho(M)<1$, it suffices to show that (I) $M$ is row substochastic, (II) for each $i^{th}$ row in $M$ with row-sum is $1$ there exists a $j^{th}$ row with row-sum less than $1$ such that there is a path from $j$ to $i$ in the associated digraph $\mathcal{G}(M)$. Then, by Theorem 4.11 in [20])1, $\rho(M)<1$.
Step-I: $M$ is row-substochastic if: (i) a node $i\in\alpha^{c}$ whose in-neighbour is $p,q$ or an oblivious agent in iSCC satisfies $\sum_{j\in\alpha^{c}}w_{ij}<1$, (ii) node $i\in\alpha^{c}$ is a stubborn agent and the corresponding row-sum is at most $(1-\beta_{i})<1$. In an arbitrary digraph $\mathcal{G}$, the following scenarios may occur:
  • $h\in\alpha^{c}$ has an oblivious agent: Due to weak connectivity of $\mathcal{G}$, an oblivious agent in iSCC exists in $\alpha$ that has a path to $h$, implying that (i) holds.
  • All agents in $\alpha^{c}$ are non-oblivious and a stubborn agent exists in $\alpha^{c}$: in this case, (ii) holds for the stubborn agent.
  • All agents in $\alpha^{c}$ are non-oblivious and non-stubborn: Since $q$ cannot be stubborn, node $p$ is the only stubborn agent in $\mathcal{G}$. Thus, at least one node exists in $\alpha^{c}$ that has either $p$ or $q$ as in-neighbour, implying that (i) holds.
Thus, there always exists a node in $\alpha^{c}$ that satisfies either (i) or (ii). Hence, $M$ is row substochastic.
Step-II: As discussed in Step-I, each oblivious agent has a path from an oblivious agent in iSCC in $\mathcal{G}$. It follows from (i) that this path traverses an oblivious agent in $\alpha^{c}$ whose row-sum in $M$ is less than $1$. A special case could occur if each path in $\mathcal{G}$ from the oblivious agents in iSCC to an oblivious agent $\tilde{j}\in\alpha^{c}$ either traverses $p$ or $q$. In this case, the in-neighbour(s) of $p$ or $q$ in $\alpha^{c}$ who has row-sum in $M$ less than $1$ lie on this path. Hence, every oblivious agent in $\alpha^{c}$ with row sum $1$ has a path in $\mathcal{G}(M)$ from a node with row sum <1<1. A similar argument applies to the non-oblivious agents as well. Thus, statement II also holds and $\rho(M)<1$.
Proof of Theorem-1: Let $q\in\mathcal{N}_{p}$. To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn. (2) using the Kron Reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (3). Let the first $u$ agents in $\mathcal{G}$ be oblivious agents that form the iSCCs. The oblivious agents in each iSCC in $\mathcal{G}$ have consensus [9]. Hence, we consider $q$ is not an oblivious agent in an iSCC; otherwise, $p$ and $q$ belong to an iSCC that is known to achieve consensus.
Consider $\alpha$ to be $\{p,q,1,...,u,n+1,...,n+m\}$. From Remark 4.3, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined. Next, we evaluate $R/\alpha^{c}$ given by eqn. (4). The following structural property of the associated network $\mathcal{G}(R)$ will be used in determining $R/\alpha^{c}$:
@@ 1. P1:
Property P1 implies that $r_{q,j}=0$ for all $j\in\{[u],n+1,...,n+m\}$ because node $j$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$. Additionally, the nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources. Thus, the corresponding rows of $R$ have all entries as zero. Therefore, $R[\alpha]$ takes the following form:
$R[\alpha]={\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}}$.
Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 5.2, it follows that
@@
Let $F:[{u}+m+2]\to\alpha$ be a mapping such that $F(1)=p,F(2)=q$, $F(j+2)=j$ for $j\in[u]$, and $F(u+2+j)=n+j$ for $j\in[m]$. This mapping relates the row (and column) indices of $Y$ (and $R/\alpha^{c}$) with the agents in set $\alpha$. Then, from eqn. (6), the entry $y_{ij}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $F(j)$ to $F(i)$ such that each node on this walk (except $F(i)$ and $F(j)$) belongs to $\alpha^{c}$. By property P1), $y_{2j}=0$ for all $j\in{\{3,...,u+m+2\}}$ because each walk from corresponding $F(j)$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$. Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the form:
$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&R_{1}&r_{1,u+2}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}
where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+2]$. The matrix $R_{1}$ is the submatrix $R/\alpha^{c}[1,[u]]$ highlighting the corresponding non-zero entries. From Lemma 5.1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0. Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$. Hence, by eliminating the states $\alpha^{c}$ in the steady state eqns. (2), we get $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$. $\blacksquare$
5.2 Proof of Corollary 2
We classified the agents in $\mathcal{G}$ into sets $\mathcal{O}$ and $\mathcal{V}\setminus\mathcal{O}$. As $\mathcal{V}\setminus\mathcal{O}$ is union of tkt-k disjoint sets $\{i\}\cup\mathcal{N}_{i}$, where each $i\in\mathcal{L}$ is an LTP agent. Hence, ktk-t LTP agents are present in set $\mathcal{V}\setminus\mathcal{O}$. By Theorem 1, each LTP agent $i$ and the agents in $\mathcal{N}_{i}$ form an opinion cluster. Therefore, the agents in set $\mathcal{V}\setminus\mathcal{O}$ form ktk-t clusters. The set $\mathcal{V}\setminus\mathcal{O}$ contains the $m$ stubborn agents. Since stubborn agents persuaded by an LTP agent, each stubborn agent must be an LTP agent. Thus, $k-t\geq m$.
[…]
Let $C_{j}$ be the set of oblivious agents in $j^{th}$ iSCC and those oblivious agents only have a path from the $j^{th}$ iSCC. The opinions of agents in $C_{j}$ evolve as $\mathbf{y}_{j}(k)=W_{i}^{j}\mathbf{y}_{j}(k)$ with $\mathbf{y}_{j}$ is the vector that denotes the opinion of agents in this set. Since the associated digraph $G(W_{i}^{t})$ is a subgraph of $\mathcal{G}$ containing a unique aperiodic iSCC, the opinions of the agents converge to consensus (Theorem 5.1 in [20])1.
Hence, the $t$ subgroups, each composed of an iSCC of oblivious agents and those having a path only from the iSCC, form $t$ opinion clusters. Collectively, the oblivious agents in $\mathcal{O}$ and the agents in $\mathcal{V}\setminus\mathcal{O}$ form $k$ opinion clusters such that $k\geq t+m$.

### v3

@@ VII Appendix
Proof:
By definition of $\alpha$, we get $\alpha^{c}\subseteq[n]$. From the construction of $R$, it follows that $R[\alpha^{c}]=I-(I-\beta[\alpha^{c}])W[\alpha^{c}]$. Let $M=(I-\beta[\alpha^{c}])W[\alpha^{c}]$. Since $W$ is row-stochastic, $M$ satisfies $M\mathbb{1}_{n}=(I-\beta)\mathbb{1}_{n}\leq\mathbb{1}_{n}$. Hence, the row-sum of any row in $M$ can be at most $1$. By the Gersgorin disk theorem, the spectrum of $M$ is a subset of a unit disk in the complex plane (Theorem 2.9 in [17]).
Further, since $M$ is a non-negative matrix, by the Perron-Frobenius Theorem, there exists a real eigenvalue $\lambda_{M}$ such that $\lambda_{M}\geq|\mu|\geq 0$, where $\mu$ are the remaining eigenvalues of $M$ (Theorem 2.12 in [17]). Thus, $\lambda_{M}\in[0,1]$.
Further, as discussed in Remark 2, the Schur complement is well defined for any $\alpha$ containing $\{n+1,....,n+m\}$ and the oblivious agents in iSCCs of $\mathcal{G}$. Thus, $R[\alpha^{c}]$ is invertible. This means that none of the eigenvalues of $M$ is equal to $1$. Thus, $\lambda_{M}<1$ and eqn. (5) follows by Neumann series. ∎
Proof of Theorem 1: Let $q\in\mathcal{N}_{p}$. To prove that the final opinion $x_{p}^{*}=x_{q}^{*}$, we reduce the matrix $R$ in eqn. (3) using the Kron Reduction and examine the relation between $x_{p}^{*}$ and $x_{q}^{*}$ in the reduced set of linear equations (4). Let the first $u$ agents in $\mathcal{G}$ be oblivious agents that form the iSCCs. The oblivious agents in each iSCC in $\mathcal{G}$ have consensus [7]. Hence, we consider $q$ is not an oblivious agent in an iSCC; otherwise, $p$ and $q$ belong to an iSCC that is known to achieve consensus.
Consider $\alpha$ to be $\{p,q,1,...,u,n+1,...,n+m\}$. From Remark 2, we know that the Kron Reduction $R/\alpha^{c}$ is well-defined. Next, we evaluate $R/\alpha^{c}$ given by eqn. (1). The following structural property of the associated network $\mathcal{G}(R)$ will be used in determining $R/\alpha^{c}$:
@@ 1. P1:
Property P1 implies that $r_{q,j}=0$ for all $j\in\{[u],n+1,...,n+m\}$ because node $j$ cannot be an in-neighbour of $q$ in $\mathcal{G}(R)$. Additionally, the nodes n+1,,n+mn+1,...,n+m in $\mathcal{G}(R)$ are sources. Thus, the corresponding rows of $R$ have all entries as zero. Therefore, $R[\alpha]$ takes the following form: $\begin{bmatrix}r_{p,p}&r_{p,q}&R[p,[u]]&r_{p,n+1}&...&r_{p,n+m}\\ r_{q,p}&r_{q,q}&\mathbb{0}&0&...&0&\\ \mathbb{0}&\mathbb{0}&R[[u]]&\mathbb{0}&\cdots&\mathbb{0}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$}.
Next, we consider $Y=R[\alpha,\alpha^{c}](R[\alpha^{c}])^{-1}R[\alpha^{c},\alpha]$. From Lemma 4, it follows that
@@
Let $F:\alpha\to[{u}+m+2]$ be a mapping such that $F(p)=1,F(q)=2$, $F(j)=j+2$ for $j\in[u]$, and $F(n+j)=u+2+j$ for $j\in[m]$. This mapping relates the agents in set $\alpha$ with the row (and column) indices of $Y$ (and $R/\alpha^{c}$). Then, from eqn. (6), the entry $y_{F(i)F(j)}$ of $Y$ is non-zero only if there exists a walk in $\mathcal{G}(R)$ from $j$ to $i$ such that each node on this walk (except $i$ and $j$) belongs to $\alpha^{c}$. By property P1), $y_{2F(j)}=0$ for all $j\in[u]\cup\{n+1,...,n+m\}$ because each walk from corresponding $j$ to $q$ in $\mathcal{G}(R)$ always passes through $p\in\alpha$. Consequently, $R/\alpha^{c}=R[\alpha]-Y$ has the form:
$\begin{bmatrix}r_{1,1}^{1}&r_{1,2}^{1}&\Delta_{1}&r_{1,u+3}^{1}&...&r_{1,u+m+2}^{1}\\ r_{2,1}^{1}&r_{2,2}^{1}&\mathbb{0}_{u\times 1}&0&...&0\\ \mathbb{0}_{u\times 1}&\mathbb{0}_{u\times 1}&R[[u]]&\mathbb{0}_{u\times 1}&\cdots&\mathbb{0}_{u\times 1}\\ \mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&\mathbb{0}_{m\times 1}&...&\mathbb{0}_{m\times 1}\end{bmatrix}$
where $r_{i,j}^{1}$ gives the entry of matrix $R/\alpha^{c}$ for $i,j\in[m+u+2]$. The vector $\Delta_{1}=[r_{1,3}^{1},r_{1,4}^{1},...,r_{1,u+2}^{1}]$ highlights the corresponding non-zero entries. From Lemma 1, we know that $R/\alpha^{c}$ is also a Laplacian matrix and the row-sum of each of its rows equals 0. Since the second row of $R/\alpha^{c}$ has only two non-zero entries, thus, $r_{2,1}^{1}=-r_{2,2}^{1}$. Moreover, by eliminating the states $\alpha^{c}$ in the steady state eqns. (3), we get $x_{p}^{*}=x_{q}^{*}$. Since $q$ is any node in $\mathcal{N}_{p}$, this holds for all nodes in $\mathcal{N}_{p}$. $\blacksquare$.